OpenStudy (shamil98):

@Luigi0210 What is the smallest possible sum of squares of two numbers if the product is -25?

OpenStudy (shamil98):

Don't answer, let luigi try :D

OpenStudy (bibby):

oh god so sorry

OpenStudy (doggy):

-slaps bibby- bad

OpenStudy (shamil98):

so luigi, where do u start?

OpenStudy (shamil98):

I said don't answer, leave it for luigi -.-

OpenStudy (bibby):

this is a terrible question shamil

OpenStudy (shamil98):

it's relatively easy.

OpenStudy (bibby):

terribly worded.

OpenStudy (anonymous):

the product of two numbers is -25?

OpenStudy (luigi0210):

I'm only 12 man, give me a break.

OpenStudy (anonymous):

LOL he's 12 guys

OpenStudy (shamil98):

\[\huge xy = -25\] \[\huge y = -25/x\] \[\huge S = x^2 + y^2\] \[\huge s(x) = x^2 + (\frac{ -25 }{ x })^2\] minimize s(x)

OpenStudy (anonymous):

what's up with people trying to insult one's intelligence here?

OpenStudy (shamil98):

i'm not, more of a learning experience..

OpenStudy (luigi0210):

Oh I see.

OpenStudy (shamil98):

so luigi, what do you get when you minimize s(x) ? i'll do another step for you \[\huge s(x) = x^2 + 625x^{-2}\] take the first derivative and solve for x.

OpenStudy (anonymous):

he claimed to be 12. You can't expect him to know calculus just yet

OpenStudy (shamil98):

he's not 12, that was a joke, are you that gullible LOL

OpenStudy (shamil98):

either that or it's sarcasm which i didnt catch on to

OpenStudy (bibby):

I still don't get how you're implying that a square can ever be negative

OpenStudy (anonymous):

it's the internet. Then again how do i know for sure?

OpenStudy (luigi0210):

\[\LARGE s'(x)=2x+\frac{625}{3 (x^3)}\] Let's just go along for the ride.. and yea, I'm 13, gosh.

OpenStudy (anonymous):

*minus*

OpenStudy (shamil98):

uh luigi dat's wrong \[\huge s'(x) = 2x -1250x^{-3}\]

OpenStudy (luigi0210):

Oh, fml -.- Why do I have to be taught at night? xD

OpenStudy (shamil98):

lol, now let's set the equation to equal zero and solve for x.

OpenStudy (doggy):

because if you don't answer, we will all make scary noises at you from your closet while your trying to sleep O:

OpenStudy (anonymous):

x = pie!!

OpenStudy (bibby):

I still don't get how you're implying that a square can ever be negative

OpenStudy (isaiah.feynman):

@sourwing its pi. lol

OpenStudy (luigi0210):

Okay, take 5 everyone.

OpenStudy (shamil98):

\[\huge 2x - 1250x^{-3} = 0\] \[\huge 2x = 1250x^{-3}\] multiply both sides by x^3 \[\huge x^3(2x) = 1250x^{-3}(x^3)\] \[\huge 2x^4 = 1250\] \[\huge x^4 = 625\] and....

OpenStudy (anonymous):

I like pie better

OpenStudy (luigi0210):

Tis 5 ლ(■_■ლ)

OpenStudy (shamil98):

plus or minus 5

OpenStudy (shamil98):

but yes.

OpenStudy (shamil98):

\[\huge S = x^2 + y^2\] \[\huge y = -25/x\] \[\huge y = -25/5 = -5\] \[\huge x = \pm 5\] \[\huge S = (5)^2 + (-5)^2\] \[\huge S = 50\]

OpenStudy (shamil98):

Now you dont HAVE to use calculus for this, i just wanted to show how it would be done lui

OpenStudy (luigi0210):

;-;

OpenStudy (bibby):

You write really horrible vague questions. this is a joke of a question. f u

OpenStudy (shamil98):

LOL, my apologies master bibby.

OpenStudy (bibby):

Thank you for finally acknowledging me good day sirs